1) Justin has a bird feeder which gets visited by an average of 18 birds every 2 hours during daylight hours. What is the probability that the bird feeder will be visited by more than 5 birds in a 40 minute period during daylight hours? (Round your answer to three decimal places.)

2) A soda bottling company’s manufacturing process is calibrated so that 99% of bottles are filled to within specifications, while 1% is not within specification. Every hour, 12 random bottles are taken from the assembly line and tested. If 2 or more bottles in the sample are not within specification, the assembly line is shut down for recalibration. What is the probability that the assembly line will be shut down, given that it is actually calibrated correctly? Use Excel to find the probability.
Answer #1

We will use Poison distribution.

We are give that Expected number of bird visited in 40 Minutes =18*40/120=6

P(More than 5 birds visit)=1-P(X<=5)

=1-0.45====>>> using excel function '''=POISSON.DIST(5,6,TRUE)''

=0.55

Hope this will be helpful. Thanks and God Bless You :)

> ^0.554

Chasity Culton Sun, Nov 21, 2021 7:57 PM